# The Economics of Information Pollution in the Age of AI: A General Equilibrium Approach to Welfare, Measurement, and Policy

### v1

@@ Introduction
Core Assumption: Technological Asymmetry.
@@
Assumption 1 (Technological Asymmetry).
@@
Proposition 1 (Cost Asymmetry).
@@
The Platform-Producer-Consumer Game
@@
Stage I: Platform Strategy Selection.
@@
Stage II: Producer Supply Decision.
@@
Stage III: Consumer Verification Decision.
@@
Equilibrium and The Threefold Market Failure
Definition 1 (Polluted Information Equilibrium).
A Subgame Perfect Nash Equilibrium (SPNE) is a strategy profile and outcome combination $\{(m^{*},\bm{\gamma}^{*}),(Q_{H}^{*},Q_{L}^{*}),V^{*},\pi^{*}\}$ such that the decisions of the platform, producers, and consumers are mutually optimal, and beliefs are consistent with outcomes.
Theorem 1 (Existence of Equilibrium).
@@
(For a proof sketch, see Appendix A.2).
Theorem 2 (Market Failure).
@@
Proposition 2 (Paradox of AI Progress).
@@
The Information Pollution Index
Based on the market equilibrium characteristics revealed by the preceding theoretical model, we construct an Information Pollution Index (IPI) that is endogenous to the social welfare function. This section aims to build a solid bridge from abstract theory to observable measurement, ensuring the index possesses both a rigorous theoretical foundation and practical applicability for policy.
Theoretical Foundation and Axiomatic Properties of the IPI
Theoretical Construction.
Definition 2 (Information Pollution Index).
@@
Property 1 (Welfare Monotonicity).
@@
Property 3 (Policy Sensitivity).
@@
Precise Definitions and Economic Interpretations of the Four Dimensions
@@
When the capability of generation technology (”offense”) far outstrips that of detection technology (”defense”), the marginal cost of producing ”lemons” falls, systemically exacerbating the other three dimensions of harm. Graphical framework reference picture 1.
@@
From Theory to Measurement: The Core Proxy Indicator System
@@
Composite Index Construction and Application
@@
The Static Optimal Policy Portfolio: A Theoretical Benchmark
@@
$\max_{m,\bm{\gamma}}\;(1-\alpha)\Pi_{P}+\alpha\bigl{[}v(Q_{H}^{\prime})-d(Q_{L}^{\prime})\bigr{]}$ (19)
@@
Theorem 3 (First-Best Policy Portfolio).
@@
Dynamic Challenge: The Lucas Critique and Policy-Induced Innovation
@@
Proposition 3 (Policy-Induced Innovation Bias).
@@
Deep Uncertainty Challenge: Knightian Uncertainty and Robust Decision-Making
@@
Proposition 4 (Optimality of the Precautionary Principle).
@@
A Synthesis: IPI-Centered Adaptive and Robust Governance
@@
Experimental Validation
To validate our theoretical framework and the efficacy of the Information Pollution Index (IPI), we designed a comprehensive suite of experiments using an Agent-Based Model (ABM). By constructing a simulation environment that captures the three-party game among content producers, consumers, and a platform, we test our core theoretical predictions and assess the reliability of the IPI as a metric for information ecosystem health.
Experimental Design and Methodology
Simulation Model Architecture.
@@
Experiment 1: Comprehensive IPI Validation
@@
Experiment 2: Theoretical Validation and Policy Analysis
Validation of the AI Progress Paradox.
@@
Analysis of Policy Interventions.
@@ Discussion
Appendix A Appendix A: AI pollution theory framework.
Figure 4: AI pollution theory framework.
Appendix B Appendix B: Model and Experimental Details
@@
B.1 Agent-Based Model Specification
@@
Producer Agent Logic.
@@
Consumer Agent Logic.
@@
Platform Agent Logic.
The platform agent adaptively adjusts its algorithmic amplification vector $\bm{\gamma}$ and moderation intensity $m$ based on market feedback. The rules follow a gradient-ascent logic, balancing revenue maximization with a user trust (retention) constraint:
@@
B.2 Model and Experimental Parameters
@@
B.3 Overview of Experimental Procedures
@@
B.4 Summary of Main Experimental Results
@@
B.7 Code and Data Availability
@@
Appendix C Appendix C: Technical Details and Proofs
@@
$\bm{\gamma}=(\gamma_{H},\gamma_{L})$ Algorithmic amplification weights
@@
C.1 Derivation of Asymmetric Production Costs
Derivation of the CES Cost Function.
@@ C.1 Derivation of Asymmetric Production Costs
Comparison of Logarithmic Elasticities.
@@
C.2 Proof Sketch for Existence of Equilibrium
@@
    Producer Stage: Given producer heterogeneity, the aggregate supply function $Q_{j}^{S}(\gamma_{H},\gamma_{L})=\int q_{j,i}^{*}(\bm{\gamma})di$ is continuous in $\bm{\gamma}$ by the Theorem of the Maximum.
@@
    Platform Stage: The platform maximizes a continuous profit function $\Pi_{P}(m,\bm{\gamma})$ over a compact strategy space $\mathcal{S}=[0,1]\times[0,\bar{\gamma}]^{2}$. By the Weierstrass Extreme Value Theorem, a maximum exists.
@@
C.3 Proof Sketches for Inefficiency and Comparative Statics
Proof of Theorem 2 (Threefold Market Failure).
@@ C.3 Proof Sketches for Inefficiency and Comparative Statics
Proof of Proposition 1 (Paradox of AI Progress).

### v2

@@ 3.3 Production Technology: The Dynamics of Pollution
3.3.2 Core Assumption: Technological Asymmetry.
@@
Assumption 3.1 (Technological Asymmetry).
@@
Proposition 3.2 (Cost Asymmetry).
@@
3.4 The Platform-Producer-Consumer Game
@@
3.4.1 Stage I: Platform Strategy Selection.
@@
3.4.2 Stage II: Producer Supply Decision.
@@
3.4.3 Stage III: Consumer Verification Decision.
@@
3.5 Equilibrium and The Threefold Market Failure
Definition 3.3 (Polluted Information Equilibrium).
A Subgame Perfect Nash Equilibrium (SPNE) is a strategy profile and outcome combination $\{(m^{*},\boldsymbol{\gamma}^{*}),(Q_{H}^{*},Q_{L}^{*}),V^{*},\pi^{*}\}$ such that the decisions of the platform, producers, and consumers are mutually optimal, and beliefs are consistent with outcomes.
Theorem 3.4 (Existence of Equilibrium).
@@ 3.5 Equilibrium and The Threefold Market Failure
Theorem 3.5 (Market Failure).
@@
Proposition 3.6 (Paradox of AI Progress).
@@
4 The Information Pollution Index and Policy Portfolio
This section develops the Information Pollution Index (IPI) as a theoretical construct derived directly from our welfare framework. Given that the precise, nuanced components of this index (e.g., welfare deadweight loss, trust decay dynamics) are not readily available as high-frequency empirical data, the IPI is designed primarily as an operational tool for our policy simulation and analysis in Section 6. It provides a welfare-linked ”dashboard” to evaluate policy interventions within a controlled computational environment. This contrasts with the goal of Section 5, which uses a more general, observable proxy for information quality to empirically test the underlying mechanisms of the commons tragedy, rather than to measure the IPI itself. Based on the market equilibrium characteristics revealed by the preceding theoretical model, we construct an Information Pollution Index (IPI) that is endogenous to the social welfare function. This section aims to build a solid bridge from abstract theory to observable measurement, ensuring the index possesses both a rigorous theoretical foundation and practical applicability for policy.
4.1 Theoretical Foundation and Axiomatic Properties of the IPI
4.1.1 Theoretical Construction.
Definition 4.1 (Information Pollution Index).
@@ 4.1 Theoretical Foundation and Axiomatic Properties of the IPI
Property 4.2 (Welfare Monotonicity).
@@
Property 4.4 (Policy Sensitivity).
@@
4.2 Precise Definitions and Economic Interpretations of the Four Dimensions
@@
When the capability of generation technology (”offense”) far outstrips that of detection technology (”defense”), the marginal cost of producing ”lemons” falls, systemically exacerbating the other three dimensions of harm. Graphical framework reference picture 3.
@@
4.3 From Theory to Measurement: The Core Proxy Indicator System
@@
4.4 Composite Index Construction and Application
@@
4.5 The Static Optimal Policy Portfolio: A Theoretical Benchmark
@@
$\max_{m,\boldsymbol{\gamma}}\;(1-\alpha)\Pi_{P}+\alpha\bigl[v(Q_{H}^{\prime})-d(Q_{L}^{\prime})\bigr]$ (21)
@@
Theorem 4.5 (First-Best Policy Portfolio).
@@
4.6 Dynamic Challenge: The Lucas Critique and Policy-Induced Innovation
@@
Proposition 4.6 (Policy-Induced Innovation Bias).
@@
4.7 Deep Uncertainty Challenge: Knightian Uncertainty and Robust Decision-Making
@@
Proposition 4.7 (Optimality of the Precautionary Principle).
@@
4.8 A Synthesis: IPI-Centered Adaptive and Robust Governance
@@
Given the multidimensional complexity of the Information Pollution Index (IPI), high-frequency empirical data is not yet available. Therefore, in Section 6, we employ computational simulations to evaluate the effectiveness of IPI as a policy instrument. Before proceeding to simulation, however, Section 5 provides empirical support for the core premise of this study—the tragedy of the digital commons”—through broader macro-level proxy variables.
5 IPI Validation and Policy Analysis in a Computational Laboratory
While our general equilibrium framework in Section 3 successfully identifies the existence and inefficiency of the static Polluted Information Equilibrium, it is, by design, silent on the dynamic pathways of how this equilibrium forms, shifts, and reacts to policy. To bridge this critical gap from static theory to dynamic application, we develop an Agent-Based Model (ABM).
This ABM serves as a computational laboratory, indispensably required for three reasons. First, it allows us to move beyond a representative agent and explicitly model heterogeneous agent interactions—capturing diverse producer strategies and consumer verification thresholds. Second, it can simulate the dynamic feedback loops our static model cannot, such as how rising pollution erodes trust, which in turn alters platform incentives and consumer behavior over time. Third, it enables the study of non-equilibrium paths, revealing how the ecosystem responds to shocks (like new AI or policies) in real-time, rather than only comparing pre- and post-shock steady states.
This rich, dynamic environment is therefore essential for rigorously testing the two primary practical contributions of this paper: the IPI’s utility as a real-time ecosystem metric and the adaptive effectiveness of our proposed policy portfolio.
5.1 Experimental Design and Methodology
5.1.1 Simulation Model Architecture.
@@ 5.1 Experimental Design and Methodology
5.2 Experiment 1: Comprehensive IPI Validation
@@
5.3 Experiment 2: Theoretical Validation and Policy Analysis
5.3.1 Validation of the AI Progress Paradox.
@@ 5.3 Experiment 2: Theoretical Validation and Policy Analysis
5.3.2 Analysis of Policy Interventions.
@@ 5.4 Discussion
6 Conclusion
6.1 Core Findings and Theoretical Contribution
This paper has revealed the mechanism by which generative AI impacts information markets: technological progress, by asymmetrically altering the production costs of different quality content, can paradoxically reduce social welfare. This Paradox of AI Progress challenges optimistic, techno-deterministic forecasts, demonstrating that without proper market design, technological innovation may exacerbate, rather than mitigate, market failures.
Our theoretical framework identifies a triad of interacting market failures—a production externality, a platform governance failure, and a trust commons externality. These forces collectively sustain an inefficient Polluted Information Equilibrium. This diagnosis moves beyond simple ”content moderation” proposals to address deeper, structural issues of economic mechanism design.
6.2 Policy Implications and Practical Value
Our research demonstrates that tackling information pollution requires a multi-instrument policy portfolio rather than reliance on a single tool. This includes short-term measures, such as implementing a real-time monitoring system based on the Information Pollution Index (IPI) to serve as a ”dashboard” for policy response. Mid-term reforms should focus on advancing legislation for platform fiduciary duties to align algorithmic recommendation with social welfare. The long-term strategy must involve building content provenance infrastructure to enhance verification efficiency at a technical level.
It is particularly important that this policy design be adaptive. As AI technology evolves rapidly, fixed rules will quickly become obsolete. The IPI provides an anchor for dynamic adjustment, enabling a ”counter-cyclical” regulatory approach: interventions tighten automatically when the index worsens and relax as it improves.
6.3 Limitations and Future Directions
This study is subject to several limitations which open avenues for future research. Theoretically, our binary quality classification simplifies a reality that is a continuous spectrum; the static model does not fully capture dynamic evolutionary processes; and we do not consider inter-platform competition or user multi-homing. On the empirical side, estimating the core parameters ($\sigma_{L}$ and $\sigma_{H}$) requires more granular industrial data, the practical application of the IPI must navigate the balance between data access and privacy protection, and evaluating policy effectiveness demands long-term longitudinal studies.
Future research should explore several key directions. These include endogenizing technological progress to investigate how policy might influence AI R&D trajectories; studying regulatory arbitrage and international cooperation mechanisms for cross-border information flows; expanding the consumer model by incorporating bounded rationality and social learning from behavioral economics; and conducting in-depth analyses of key sectors such as journalism, education, and healthcare.
6.4 Concluding Remarks
Appendix A Appendix A: Model and Experimental Details
@@
A.1 B.1 Agent-Based Model Specification
@@
A.1.1 Producer Agent Logic.
@@
A.1.2 Consumer Agent Logic.
@@
A.1.3 Platform Agent Logic.
The platform agent adaptively adjusts its algorithmic amplification vector $\boldsymbol{\gamma}$ and moderation intensity $m$ based on market feedback. The rules follow a gradient-ascent logic, balancing revenue maximization with a user trust (retention) constraint:
@@
A.2 B.2 Model and Experimental Parameters
@@
A.3 B.3 Overview of Experimental Procedures
@@
A.4 B.4 Summary of Main Experimental Results
@@
A.5 B.7 Code and Data Availability
@@
Appendix B Appendix C: Technical Details and Proofs
@@
$\boldsymbol{\gamma}=(\gamma_{H},\gamma_{L})$ Algorithmic amplification weights
@@
B.1 C.1 Derivation of Asymmetric Production Costs
B.1.1 Derivation of the CES Cost Function.
@@
B.1.2 Comparison of Logarithmic Elasticities.
@@
B.2 C.2 Proof Sketch for Existence of Equilibrium
@@
    Producer Stage: Given producer heterogeneity, the aggregate supply function $Q_{j}^{S}(\gamma_{H},\gamma_{L})=\int q_{j,i}^{*}(\boldsymbol{\gamma})di$ is continuous in $\boldsymbol{\gamma}$ by the Theorem of the Maximum.
@@
    Platform Stage: The platform maximizes a continuous profit function $\Pi_{P}(m,\boldsymbol{\gamma})$ over a compact strategy space $\mathcal{S}=[0,1]\times[0,\bar{\gamma}]^{2}$. By the Weierstrass Extreme Value Theorem, a maximum exists.
@@
B.3 C.3 Proof Sketches for Inefficiency and Comparative Statics
B.3.1 Proof of Theorem 2 (Threefold Market Failure).
@@
B.3.2 Proof of Proposition 1 (Paradox of AI Progress).